# Stability of weighted minimal hypersurfaces under a lower $1$-weighted Ricci curvature bound

### v1

@@ A Lorentzian manifold $(L,\mathfrak{g})$ is called a static model if it can be written in the form of
for some Riemannian manifold $(M,\mathsf{g})$ and positive function $V\in C^{\infty}(M)$. Under the Einstein equation, a static model $(L,\mathfrak{g})$ in the form of (1.3) satisfies NEC (i.e., the stress-energy tensor is non-negative for every null vector) if and only if a triple $(M,\mathsf{g},V)$ satisfies the following substatic condition (see [55, Lemma 3.8], [5, Appendix A.1]):
@@ 1.3. Aims
for $K\in\mathbb{R}$ (see [4], [8], [9], [29], [30], [31], [37], [40] and references therein). The study under this setting may be contributed to that of the so-called self-shrinker, which is a self-similar solution to the mean curvature flow, and also can be viewed as an $f$-minimal hypersurface in the Gaussian space. We will study our object under a curvature condition
@@
for $\kappa\in\mathbb{R}$, which was introduced by Wylie [56] and Wylie-Yeroshkin [57] from the viewpoint of the study of affine connections (see also the works by Kuwae-Li [36] and Lu-Minguzzi-Ohta [41] for the interpolation between the curvature-dimension condition and (1.6)). They developed comparison geometry under the curvature bound (1.6), and $f$-minimal hypersurfaces could be found in model spaces (see Remarks 2.5 and 2.8 below). Notice that the curvature bound (1.6) with $\kappa=0$ is weaker than the classical one (1.5) with $K=0$ by the monotonicity (1.2). We extend several previous results under the curvature bound (1.5) to our setting. As explained in the previous subsection, our results may be contributed to Lorentzian extrinsic geometric analysis.
The present paper is organized as follows: In Section 2, we review basics of stability of $f$-minimal hypersurfaces (see Subsection 2.1). We also present a key ingredient of the proof concerning the revision of stability inequality in terms of the $1$-weighted Ricci curvature (see Lemma 2.1). We further collect comparison geometric results under the curvature bound (1.6) (see Subsection 2.2). In Section 3, we first study a Schoen-Yau type criterion for stable $f$-minimal hypersurfaces in weighted manifolds of non-negative $1$-weighted Ricci curvature (see Subsection 3.1). Having it at hand, we investigate the structure of three-dimensional weighted manifolds of non-negative $1$-weighted Ricci curvature (see Theorem 3.5). In Section 4, we provide non-existence results for stable $f$-minimal hypersurfaces in weighted manifolds under the positivity of $1$-weighted Ricci curvature (see Subsection 4.1). We apply them to conclude smooth compactness theorems (see Theorems 4.6 and 4.7).
@@ 2.1. Stability of weighted minimal hypersurfaces
An immersed hypersurface $\Sigma$ is said to be $f$-minimal if it is a critical point of the weighted volume functional. By the first variation formula (see e.g., [40, Proposition 1], [9, Proposition 2]), $\Sigma$ is $f$-minimal if and only if $\mathbf{H}_{f,\Sigma}\equiv 0$. An immersed $f$-minimal hypersurface $\Sigma$ is called stable if for every compactly supported normal variation of $\Sigma$, the second derivative of the weighted volume functional at $\Sigma$ is non-negative. Due to the second variation formula (see e.g., [40, Proposition 1], [9, Proposition 13]), $\Sigma$ is stable if and only if the index form
@@
This observation also can be found in recent works by Colombo-Freitas-Mari-Rigoli [13] and the first named author [19] (see [13, Section 3] and [19, Proposition 2.3]).
@@
for $K\in\mathbb{R}$ in the literature of finding stationary points for the volume functional on the space of constant scalar curvature metrics (see [42], [43], [14]). From this point of view, the natural curvature condition might be the following (see also [38, Corollary 1.5]):
@@ 2.2. Comparison geometry
We next recall the comparison geometry under the curvature bound (1.6) established by Wylie [56] and Wylie-Yeroshkin [57].
Wylie [56] has established the following splitting theorem of Cheeger-Gromoll type under the non-negativity of $1$-weighted Ricci curvature (see [56, Theorem 1.2]):
@@
In the splitting case in Theorem 2.4, one can observe that each slice (i.e., each level set of the Busemann function) must be an $f$-minimal hypersurface since the Busemann function is $f$-harmonic and its norm of the gradient is identically one from its proof (see also the example in [56, Corollary 2.4]).
Wylie-Yeroshkin [57] have developed Laplacian and volume comparison under the curvature bound (1.6) (see [57, Section 4]). They have derived the following:
@@
As a consequence of volume comparison, Wylie-Yeroshkin [57] have concluded the following finiteness of the fundamental group (see [57, Theorem 2.9]):
Theorem 2.7 ([57]).
Let $(M,g,f)$ be an $n$-dimensional complete weighted manifold. For $\kappa>0$, we assume $\operatorname{Ric}_{f}^{1}\geq(n-1)\,\kappa\,\mathrm{e}^{-\frac{4f}{n-1}}g$. Then the fundamental group $\pi_{1}(M)$ is finite.
@@
Under the curvature bound (1.6) with $\kappa>0$, Wylie-Yeroshkin [57] have produced a comparison theorem of Bonnet-Myers type, and a rigidity theorem of Cheng type by setting a model space as a warped product in the following form (see [57, Theorems 2.2 and 2.6]):
@@ 3.1. Schoen-Yau type criterion
In the present subsection, we present a Schoen-Yau type criterion for stable $f$-minimal hypersurfaces in weighted manifolds of non-negative $1$-weighted Ricci curvature. Such a result has been obtained by Liu [40] under the non-negativity of $\infty$-weighted Ricci curvature (see [40, Sections 3, 5]). We generalize it under a weak assumption.
We first show the following (cf. [40, Theorem 1 and Proposition 2]):
@@
In [40, Theorem 1], he has also shown a rigidity result for weighed manifold of non-negative $\infty$-weighted Ricci curvature under the existence of a certain compact $f$-area-minimizing hypersurface. Similarly, we have the following: Let $(M,g,f)$ be an $n$-dimensional complete orientable weighted manifold. We assume $\operatorname{Ric}^{1}_{f}\geq 0$. Let $\Sigma$ be a closed orientable embedded stable $f$-minimal hypersurface in $M$. If $\Sigma$ is $f$-area-minimizing in its homology class, then $M$ is isometric to a quotient of warped product in the form of
@@
with $f(\gamma_{z}(t))=f_{1}(t)+f_{2}(z)$ for some $f_{1}:\mathbb{R}\to\mathbb{R}$ and $f_{2}:\Sigma\to\mathbb{R}$, where $\gamma_{z}$ is the line with $\gamma_{z}(0)=z$ that is orthogonal to $\Sigma$. The outline of the proof is as follows: Let $\{\Sigma_{t}\}_{t\in(-\varepsilon,\varepsilon)}$ be a family of hypersurfaces in $M$ such that $\Sigma_{t}$ is the $t$-level set of the signed distance function from $\Sigma$ in $M$. On the side such that $t\geq 0$, the weighted Laplacian of the distance function from $\Sigma$ is non-positive due to the Laplacian comparison (see e.g., [56, Theorem 5.2], [50, Lemmas 3.3 and 3.8]); in particular, the weighted measure of $\Sigma_{t}$ is non-increasing in $t$. Since $\Sigma=\Sigma_{0}$ is $f$-area-minimizing, it must be constant in $t$, and hence the equality in the Laplacian comparison holds, and the rigidity for the equality case can be used. By applying the same argument to the opposite side, and by induction, we arrive at the desired assertion thanks to [56, Proposition 2.2].
One can conclude the following (cf. [40, Proposition 3]):
@@ 3.2. Structure of weighted three-manifolds
The structure of complete three-manifolds of non-negative Ricci curvature has been investigated by Hamilton [24], [25], Schoen-Yau [52] and Liu [39]. As a consequence, Liu [39] has concluded that the Milnor conjecture is true in the three dimensional case (see [39, Corollary 1]). In [40], he has also examined the structure of three-dimensional weighted manifolds of non-negative $\infty$-weighted Ricci curvature via the analysis of stable $f$-minimal hypersurfaces (see [40, Theorem 3]). We generalize his result for weighted manifolds of non-negative $1$-weighted Ricci curvature.
@@ 3. $(2c)$
We first consider the case where $M$ is compact. If the universal cover of $M$ is compact, then it is diffeomorphic to $\mathbb{S}^{3}$ due to the resolution of the Poincaré conjecture ([47], [48], [49]). If the universal cover is non-compact, then it contains a line (see e.g., [56, Theorem 4.5]), and hence Theorem 2.4 tells us that the universal cover is isometric to a warped product.
We now study the case where $M$ is non-compact. If $\pi_{2}(M)$ is non-trivial, then the universal cover contains a line (see [52, Lemma 2]), and thus the universal cover is isometric to a warped product by Theorem 2.4. If $\pi_{2}(M)$ is trivial, then by the Hurewicz theorem, $M$ is a $K(\pi,1)$ space since $M$ is three dimensional and non-compact (see [52, Lemma 3]); in particular, if $\pi_{1}(M)$ is also trivial, then $M$ is contractible thanks to the Whitehead theorem. Therefore, it suffices to discuss the case where $\pi_{1}(M)$ is non-trivial.
We assume that $\pi_{2}(M)$ is trivial, and $\pi_{1}(M)$ is non-trivial. From the the fact that $M$ is a $K(\pi,1)$ space, it follows that $\pi_{1}(M)$ is torsion free; in particular, we may assume that $\pi_{1}(M)=\mathbb{Z}$ and $M$ is orientable by passing to a suitable covering. Let $\gamma$ be a Jordan curve that represents the generator of $\pi_{1}(M)$. By the argument of Anderson [1], there is a complete orientable immersed $f$-area-minimizing surface in $M$ that intersects with $\gamma$ (see [1, Lemma 2.2], and also [40, Lemma 2]). From Proposition 3.4, it follows that (3.3) holds along it. If $\operatorname{Ric}^{1}_{f}>0$ on $M$, then this is a contradiction.
@@
if $t\leq R^{-5}/2$. Hence, similarly to [40], one can conclude that there exists a sufficiently small $R>0$ such that for every sufficiently small $t$, the metric $g_{t}$ has positive $1$-weighted Ricci curvature over $B_{R}(o)\backslash B_{7R/8}(o)$. We apply this perturbation finitely many times such that the $1$-weighted Ricci curvature is positive along $\gamma$, and also non-negative except for a small neighborhood of $o$. For this new metric, we construct a complete orientable immersed $f$-area-minimizing surface in $M$ that intersects with $\gamma$ by using the argument in [1] again. Notice that it also has the quadratic $f$-volume growth by Proposition 2.6 since the new metric is uniformly equivalent to the original one $g$ (cf. [40, (6.3)]). It follows that the surface must pass through the small neighborhood of $o$. Indeed, if this is not true, then the $1$-weighted Ricci curvature is non-negative along the surface, and also positive at the intersection with $\gamma$; however, this contradicts with Lemma 3.1.
@@ 2. $(2c)^{\ast}$
We first discuss $(2c)^{\ast}$. By the additional assumption, we see that for each $o\in M$, the surface $\Sigma$ is uniquely determined, and $M$ is foliated by them (cf. [54, Section 3]). Let $\{\Sigma_{t}\}_{t}$ stand for the foliation. Notice the weighted evolution formula
@@
holds for the variation field $\phi\,\nu$ on $\Sigma_{t}$, which can be derived from the evolution formula for the mean curvature and unit normal vector filed in the unweighted case (see e.g., [27, Theorem 3.2]). From the $f$-minimality, (3.5) and Lemma 2.1, it follows that
@@
By a Liouville result for $f$-harmonic functions together with (3.4), $\phi$ must be constant (see [40, Lemma 3]). By the re-parametrization, we may assume $\phi\equiv 1$, and conclude $\nabla_{\nu}\nu=0$. Notice that the second fundamental form of $\Sigma_{t}$ coincides with the Hessian of the distance function from $\Sigma_{0}$, and satisfies (3.5). Thus, by the same argument as in the proof of Theorem 2.4 in [56], the universal cover is isometric to a warped product over $\mathbb{R}$ in the form of
@@
The assumption for upper boundedness of $f$ in Theorem 2.4 has been weakened as the $f$-completeness introduced in [56], [57] (cf. [56, Remark 1.2]). In the same manner, the assumption in Theorem 3.5 also can be weakened as the combination of the $f$-completeness and the lower boundedness of $f$.

### v2

@@ A Lorentzian manifold $(L,\mathfrak{g})$ is called a static model if it can be written in the form of
for some Riemannian manifold $(M,\mathsf{g})$ and positive function $V\in C^{\infty}(M)$. Under the Einstein equation, a static model $(L,\mathfrak{g})$ in the form of (1.3) satisfies NEC (i.e., the stress-energy tensor is non-negative for every null vector) if and only if a triple $(M,\mathsf{g},V)$ satisfies the following substatic condition (see [59, Lemma 3.8], [5, Appendix A.1]):
@@ 1.3. Aims
for $K\in\mathbb{R}$ (see [4], [8], [9], [31], [32], [33], [40], [43] and references therein). The study under this setting may be contributed to that of the so-called self-shrinker, which is a self-similar solution to the mean curvature flow, and also can be viewed as an $f$-minimal hypersurface in the Gaussian space. We will study our object while keeping a curvature condition
@@
for $\kappa\in\mathbb{R}$ in mind, which was introduced by Wylie [60], Wylie-Yeroshkin [61] from the viewpoint of the study of affine connections (see also the works by Kuwae-Li [38] and Lu-Minguzzi-Ohta [44] for the interpolation between the curvature-dimension condition and (1.6)). They developed comparison geometry under the curvature bound (1.6), and $f$-minimal hypersurfaces could be found in model spaces (see Remarks 2.5 and 2.9 below). Notice that the curvature bound (1.6) with $\kappa=0$ is weaker than the classical one (1.5) with $K=0$ by the monotonicity (1.2). We extend several previous results under the curvature bound (1.5) to our setting. As explained in the previous subsection, our results may be contributed to Lorentzian extrinsic geometric analysis.
The present paper is organized as follows: In Section 2, we review basics of stability of $f$-minimal hypersurfaces (see Subsection 2.1). We also present a key ingredient of the proof concerning the revision of stability inequality in terms of the $1$-weighted Ricci curvature (see Lemma 2.1). We further collect comparison geometric results under the curvature bound (1.6) (see Subsection 2.2). In Section 3, we first study a Schoen-Yau type criterion for stable $f$-minimal hypersurfaces in weighted manifolds of non-negative $1$-weighted Ricci curvature (see Subsection 3.1). Having it at hand, we investigate the structure of three-dimensional weighted manifolds of non-negative $1$-weighted Ricci curvature (see Theorem 3.5). In Section 4, we provide non-existence results for stable $f$-minimal hypersurfaces in weighted manifolds under the positivity of $1$-weighted Ricci curvature (see Subsection 4.1). We apply them to conclude a smooth compactness theorem (see Theorem 4.4).
@@ 2.1. Stability of weighted minimal hypersurfaces
An immersed hypersurface $\Sigma$ is said to be $f$-minimal if it is a critical point of the weighted volume functional. By the first variation formula (see e.g., [43, Proposition 1], [9, Proposition 2]), $\Sigma$ is $f$-minimal if and only if $\mathbf{H}_{f,\Sigma}\equiv 0$. An immersed $f$-minimal hypersurface $\Sigma$ is called stable if for every compactly supported normal variation of $\Sigma$, the second derivative of the weighted volume functional at $\Sigma$ is non-negative. Due to the second variation formula (see e.g., [43, Proposition 1], [9, Proposition 13]), $\Sigma$ is stable if and only if the index form
@@
This observation also can be found in recent works by Colombo-Freitas-Mari-Rigoli [14] and the first named author [20] (see [14, Section 3] and [20, Proposition 2.3]).
@@
for $K\in\mathbb{R}$ in the literature of finding stationary points for the volume functional on the space of constant scalar curvature metrics (see [46], [47], [15]). From this point of view, the natural curvature condition might be the following (see also [41, Corollary 1.5]):
@@ 2.2. Comparison geometry
We next recall the comparison geometry under the curvature bound (1.6) established by Wylie [60] and Wylie-Yeroshkin [61] (see also [39]).
Wylie [60] has established the following splitting theorem of Cheeger-Gromoll type under the non-negativity of $1$-weighted Ricci curvature (see [60, Theorem 1.2]):
@@
In the splitting case in Theorem 2.4, one can observe that each slice (i.e., each level set of the Busemann function) must be an $f$-minimal hypersurface since the Busemann function is $f$-harmonic and its norm of the gradient is identically one from its proof (see also the example in [60, Corollary 2.4]).
Wylie-Yeroshkin [61] have developed Laplacian and volume comparison under the curvature bound (1.6) (see [61, Section 4]). They have derived the following:
@@
We next discuss results under the positivity of $1$-weighted Ricci curvature. Let $(M,g,f)$ be an $n$-dimensional complete weighted manifold. Wylie-Yeroshkin [61] have introduced the reparametrized distance $d_{f}:M\times M\to\mathbb{R}$ defined by
$d_{f}(x,y):=\inf_{\gamma}\int^{d(x,y)}_{0}\,\mathrm{e}^{-\frac{2f(\gamma(\xi))}{n-1}}\,\mathrm{d}\xi,$
where the infimum is taken over all unit speed minimal geodesics $\gamma:[0,d(x,y)]\to M$ from $x$ to $y$. They obtained the following estimate of Bonnet-Myers type (see [61, Theorem 2.2]):
Proposition 2.7 ([61]).
Let $(M,g,f)$ be an $n$-dimensional complete weighted manifold. For $\kappa>0$, we assume $\operatorname{Ric}_{f}^{1}\geq(n-1)\,\kappa\,\mathrm{e}^{-\frac{4f}{n-1}}g$. Then we have
$\sup_{x,y\in M}d_{f}(x,y)\leq\frac{\pi}{\sqrt{\kappa}}.$
In particular, if $f$ is bounded from above, then $M$ must be compact.
@@
In Subsection 4.2 below, we will focus on a closed weighted manifold $(M,g,f)$ of positive $1$-weighted Ricci curvature. Then it satisfies the curvature bound (1.6) for some $\kappa>0$, and the curvature bound can be lifted over the universal cover. Since the lift of $f$ is also bounded from above, Proposition 2.7 says that the universal cover must be compact; in particular, the fundamental group $\pi_{1}(M)$ must be finite. Notice that the finiteness of the fundamental group can be available under a weaker setting that $(M,g,f)$ is complete and satisfies the curvature bound (1.6) for $\kappa>0$ (see [61, Theorem 2.9]).
Remark 2.9.
Under the same setting as in Proposition 2.7, Wylie-Yeroshkin [61] have produced a rigidity theorem of Cheng type concerning the equality case by setting a model space as a warped product of the following form (see [61, Theorem 2.6]):
@@
We close this subsection with the following Frankel property (see [22, Proposition 3.6]):
Proposition 2.10 ([22]).
Let $(M,g,f)$ be a complete weighted manifold of positive $1$-weighted Ricci curvature. Let $\Sigma_{1}$ and $\Sigma_{2}$ be two complete properly immersed $f$-minimal hypersurfaces in $M$ such that at least one of them is compact. Then $\Sigma_{1}$ and $\Sigma_{2}$ must intersect.
@@ 3.1. Schoen-Yau type criterion
In the present subsection, we present a Schoen-Yau type criterion for stable $f$-minimal hypersurfaces in weighted manifolds of non-negative $1$-weighted Ricci curvature. Such a result has been obtained by Liu [43] under the non-negativity of $\infty$-weighted Ricci curvature (see [43, Sections 3, 5]). We generalize it under a weak assumption.
We first show the following (cf. [43, Theorem 1 and Proposition 2]):
@@
In [43, Theorem 1], he has also shown a rigidity result for weighed manifold of non-negative $\infty$-weighted Ricci curvature under the existence of a certain compact $f$-area-minimizing hypersurface. Similarly, we have the following: Let $(M,g,f)$ be an $n$-dimensional complete orientable weighted manifold. We assume $\operatorname{Ric}^{1}_{f}\geq 0$. Let $\Sigma$ be a closed orientable embedded stable $f$-minimal hypersurface in $M$. If $\Sigma$ is $f$-area-minimizing in its homology class, then $M$ is isometric to a quotient of warped product in the form of
@@
with $f(\gamma_{z}(t))=f_{1}(t)+f_{2}(z)$ for some $f_{1}:\mathbb{R}\to\mathbb{R}$ and $f_{2}:\Sigma\to\mathbb{R}$, where $\gamma_{z}$ is the line with $\gamma_{z}(0)=z$ that is orthogonal to $\Sigma$. The outline of the proof is as follows: Let $\{\Sigma_{t}\}_{t\in(-\varepsilon,\varepsilon)}$ be a family of hypersurfaces in $M$ such that $\Sigma_{t}$ is the $t$-level set of the signed distance function from $\Sigma$ in $M$. On the side such that $t\geq 0$, the weighted Laplacian of the distance function from $\Sigma$ is non-positive due to the Laplacian comparison (see e.g., [60, Theorem 5.2], [54, Lemmas 3.3 and 3.8]); in particular, the weighted measure of $\Sigma_{t}$ is non-increasing in $t$. Since $\Sigma=\Sigma_{0}$ is $f$-area-minimizing, it must be constant in $t$, and hence the equality in the Laplacian comparison holds, and the rigidity for the equality case can be used. By applying the same argument to the opposite side, and by induction, we arrive at the desired assertion thanks to [60, Proposition 2.2].
One can conclude the following (cf. [43, Proposition 3]):
@@ 3.2. Structure of weighted three-manifolds
The structure of complete three-manifolds of non-negative Ricci curvature has been investigated by Hamilton [26], [27], Schoen-Yau [56] and Liu [42]. As a consequence, Liu [42] has concluded that the Milnor conjecture is true in the three dimensional case (see [42, Corollary 1]). In [43], he has also examined the structure of three-dimensional weighted manifolds of non-negative $\infty$-weighted Ricci curvature via the analysis of stable $f$-minimal hypersurfaces (see [43, Theorem 3]). We generalize his result for weighted manifolds of non-negative $1$-weighted Ricci curvature.
@@ 3. $(2c)$
We first consider the case where $M$ is compact. If the universal cover of $M$ is compact, then it is diffeomorphic to $\mathbb{S}^{3}$ due to the resolution of the Poincaré conjecture ([51], [52], [53]). If the universal cover is non-compact, then it contains a line (see e.g., [60, Theorem 4.5]), and hence Theorem 2.4 tells us that the universal cover is isometric to a warped product.
We now study the case where $M$ is non-compact. If $\pi_{2}(M)$ is non-trivial, then the universal cover contains a line (see [56, Lemma 2]), and thus the universal cover is isometric to a warped product by Theorem 2.4. If $\pi_{2}(M)$ is trivial, then by the Hurewicz theorem, $M$ is a $K(\pi,1)$ space since $M$ is three dimensional and non-compact (see [56, Lemma 3]); in particular, if $\pi_{1}(M)$ is also trivial, then $M$ is contractible thanks to the Whitehead theorem. Therefore, it suffices to discuss the case where $\pi_{1}(M)$ is non-trivial.
We assume that $\pi_{2}(M)$ is trivial, and $\pi_{1}(M)$ is non-trivial. From the the fact that $M$ is a $K(\pi,1)$ space, it follows that $\pi_{1}(M)$ is torsion free; in particular, we may assume that $\pi_{1}(M)=\mathbb{Z}$ and $M$ is orientable by passing to a suitable covering. Let $\gamma$ be a Jordan curve that represents the generator of $\pi_{1}(M)$. By the argument of Anderson [1], there is a complete orientable immersed $f$-area-minimizing surface in $M$ that intersects with $\gamma$ (see [1, Lemma 2.2], and also [43, Lemma 2]). From Proposition 3.4, it follows that (3.3) holds along it. If $\operatorname{Ric}^{1}_{f}>0$ on $M$, then this is a contradiction.
@@
if $t\leq R^{-5}/2$. Hence, similarly to [43], one can conclude that there exists a sufficiently small $R>0$ such that for every sufficiently small $t$, the metric $g_{t}$ has positive $1$-weighted Ricci curvature over $B_{R}(o)\backslash B_{7R/8}(o)$. We apply this perturbation finitely many times such that the $1$-weighted Ricci curvature is positive along $\gamma$, and also non-negative except for a small neighborhood of $o$. For this new metric, we construct a complete orientable immersed $f$-area-minimizing surface in $M$ that intersects with $\gamma$ by using the argument in [1] again. Notice that it also has the quadratic $f$-volume growth by Proposition 2.6 since the new metric is uniformly equivalent to the original one $g$ (cf. [43, (6.3)]). It follows that the surface must pass through the small neighborhood of $o$. Indeed, if this is not true, then the $1$-weighted Ricci curvature is non-negative along the surface, and also positive at the intersection with $\gamma$; however, this contradicts with Lemma 3.1.
@@ 2. $(2c)^{\ast}$
We first discuss $(2c)^{\ast}$. By the additional assumption, we see that for each $o\in M$, the surface $\Sigma$ is uniquely determined, and $M$ is foliated by them (cf. [58, Section 3]). Let $\{\Sigma_{t}\}_{t}$ stand for the foliation. Notice the weighted evolution formula
@@
holds for the variation field $\phi\,\nu$ on $\Sigma_{t}$, which can be derived from the evolution formula for the mean curvature and unit normal vector filed in the unweighted case (see e.g., [29, Theorem 3.2]). From the $f$-minimality, (3.5) and Lemma 2.1, it follows that
@@
By a Liouville result for $f$-harmonic functions together with (3.4), $\phi$ must be constant (see [43, Lemma 3]). By the re-parametrization, we may assume $\phi\equiv 1$, and conclude $\nabla_{\nu}\nu=0$. Notice that the second fundamental form of $\Sigma_{t}$ coincides with the Hessian of the distance function from $\Sigma_{0}$, and satisfies (3.5). Thus, by the same argument as in the proof of Theorem 2.4 in [60], the universal cover is isometric to a warped product over $\mathbb{R}$ in the form of
@@
The assumption for upper boundedness of $f$ in Theorem 2.4 has been weakened as the $f$-completeness introduced in [60], [61] (cf. [60, Remark 1.2]). In the same manner, the assumption in Theorem 3.5 also can be weakened as the combination of the $f$-completeness and the lower boundedness of $f$.